Joe348:

the height of a trapezoid can be expressed as x-4, while the bases can be expressed as x+4 and x+9. If the area of the trapezoid is 99 cm^2 Find the length if t

snowflake0531:

I think you didn't fully copy the question over

Joe348:

one sec

Joe348:

1 attachment
snowflake0531:

right so the larger base is obviously x + 9, so you just need to find x

snowflake0531:

do you know how to find the area of a trapezoid?

Joe348:

A=1/2(b1+b2)?

snowflake0531:

yes, multiplied by height

Joe348:

so like 99=1/2(x-4+x+9)?

snowflake0531:

close, \[99=(x-4)(\frac{(x+4)+(x+9)}{2})\]

Joe348:

wait why the 2?

snowflake0531:

\[ 99 = (x-4)(\frac{2x+13}{2}) \\ 198=(x-4)(2x+13)\]

snowflake0531:

@joe348 wrote:
wait why the 2?
1/2(a+b) = (a+b)/2

Joe348:

oh okay

Joe348:

so now what?

snowflake0531:

can you solve for x?

Joe348:

wait im confused

Joe348:

what x?

snowflake0531:

@snowflake0531 wrote:
\[ 99 = (x-4)(\frac{2x+13}{2}) \\ 198=(x-4)(2x+13)\]

Joe348:

\[0=2x^2+5x-250 \] and \[0=(2x+25)(x-10)\] right?

snowflake0531:

if you're taking it from the internet... why post in the first place

Joe348:

because like i dont understand it

Joe348:

i understand what i put but not what you put

snowflake0531:

@joe348 wrote:
\[0=2x^2+5x-250 \] and \[0=(2x+25)(x-10)\] right?
you understand how they got this equation?

snowflake0531:

then what is x

Joe348:

@snowflake0531 wrote:
@snowflake0531 wrote:
\[ 99 = (x-4)(\frac{2x+13}{2}) \\ 198=(x-4)(2x+13)\]
from this i meant

Joe348:

ohh okay i got it so all i really need to do is factor right?

snowflake0531:

you already have it factored, you just need to solve for x, in what you have

Joe348:

so like x=10 right? then i plug that in to solve for A?

Joe348:

or the length?

snowflake0531:

x+9 is the longer height so just 10 + 9

Joe348:

oh so its 19 and 14

snowflake0531:

yes

Joe348:

bruh why didnt you just tell me thats all i needed to do😰

snowflake0531:

tell you what?

Joe348:

idk but i got the final answer

Joe348:

brb

snowflake0531:

good job

Joe348:

thanks for your help c.x

snowflake0531:

yw~